OpenStudy (anonymous):

The polynomial 2x^2-3x+5, also, can be called a

jimthompson5910 (jim_thompson5910):

This is a quadratic since the degree is 2 This is a trinomial since there are 3 terms So it's a quadratic trinomial

OpenStudy (saifoo.khan):

a quadratic expression.

OpenStudy (ybarrap):

since (-3)^2 - 4*(2)*(5) = 9 - 40 = - 31 < 0, this poly has complex-conjugate roots. So, this quadratic polynomial will never cross the x-axis, yet will be zero for two special (complex) numbers.